OpenStudy (anonymous):

If you have a problem like \[\frac{3}{\sqrt{-9}}\]does it matter if you bring the i out first or rationalize first? I believe that it doesn't matter which because it'll come out as i?

OpenStudy (phi):

it does not matter, but generally people write \[ \sqrt{-9} \text{ as }i\sqrt{9} \] so \[ \frac{\cancel{3}}{\cancel{3}i}= \frac{i}{i*i}= -i \]

OpenStudy (anonymous):

Wait, if you do it the other way, wouldn't it become i? \[\frac{3}{\sqrt{-9}} \times \frac{\sqrt{-9}}{\sqrt{-9}} \implies \frac{9i}{9} \implies i\]Or did I do something wrong?

OpenStudy (phi):

sqrt(-9)*sqrt(-9)= -9

OpenStudy (anonymous):

Ohh...I thought that you would multiply was was inside before simplify the radical...I guess I was wrong. THank you :)

OpenStudy (phi):

I think \[ \sqrt{a \cdot b} =\sqrt{a}\cdot \sqrt{b} \] only works for positive a,b

OpenStudy (anonymous):

Oh ok. THat makes sense since you can't really have two answers for something like this. Again, thank you. I'll double check the rule in a second.

OpenStudy (anonymous):

Yup. THat is the case. Would that also apply to dividing radicals?